OpenStudy (anonymous):

can anyone help im combining like terms and i ended up -21 which isnt a choice

OpenStudy (anonymous):

What is the value of z in the equation 3z + 5 = 4z − 8? A. −13 B. −1 C. 1 D. 13

OpenStudy (anonymous):

@teddydaddy23 @imqwerty @Bigbosssaint21

OpenStudy (bigbosssaint21):

i believe its D

OpenStudy (anonymous):

yes it is D... wanna know a simple solution?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what is 3 x 13 + 5 =

OpenStudy (anonymous):

44

OpenStudy (anonymous):

what is 4 x 13 - 8 =

OpenStudy (anonymous):

44

OpenStudy (anonymous):

ok, a simple way to do this is to just plug in each choice for the letter and see which one works

OpenStudy (anonymous):

ohh ok i see thanks so much

OpenStudy (anonymous):

or use this site ^

OpenStudy (anonymous):

just type in your problem and it will give you an answer each time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good luck in the future.. if u need help tag me!!!!

OpenStudy (anonymous):

got it thanks so much @Skielerlucas04